# The Jacobi Factoring Circuit: Quantum Factoring with Near-Linear Gates and Sublinear Space and Depth

### v1

@@ 5 Fully Factoring Integers with Distinct Exponents in their Prime Factorization
We say an integer $N$ is special if all the exponents in its prime factorization are distinct i.e. we can write $N=p_{1}^{\alpha_{1}}\ldots p_{r}^{\alpha_{r}}$ for distinct primes $p_{1},\ldots,p_{r}$ and distinct positive integers $\alpha_{1},\ldots,\alpha_{r}$.
Remark.
One might wonder whether special integers turn out to be classically easy to factor as well. We state some evidence that this is not the case here. Indeed, the density of these integers was studied by Aktaş and Murty [AM17], who showed that for any integer $N_{\mathrm{max}}$, the number of special integers $N\in[1,N_{\mathrm{max}}]$ is $\approx\frac{1.7N_{\mathrm{max}}}{\ln N_{\mathrm{max}}}$. The following classes of integers that are classically even slightly easier than general to fully factor do not contain enough elements in total to cover all special numbers:
@@
With this in mind, we now turn our attention to fully factoring special integers. We first begin with a simple lemma. At a high level, we want to show that the $\Omega(1)$ success probability in Theorem 3.1 can be boosted to be very close to 1. This is not obvious since given $N=A^{2}B$ with $B$ squarefree, it may not be possible to efficiently determine whether the algorithm has succeeded in recovering $B$. We show that this is not difficult to work around.
@@
Let $N$ be a positive integer, with unique representation as $N=A^{2}B$ for $B$ squarefree. Moreover, we say $N$ is very good if it is composite and neither squarefree nor a square.
@@
Then there exists another algorithm $\mathcal{B}$ that, given a special integer $N$ as input, recovers the full prime factorization of $N$ with probability $1-\mathsf{negl}(\log N)$. This algorithm makes at most $O(\sqrt{\log N})$ calls to $\mathcal{A}$ with inputs $N^{\prime}$ that are always $\leq N$ and with repetition parameter $T=\omega(\log N)$. Outside of calls made to $\mathcal{A}$ the algorithm is classical and runs in time $\mathsf{poly}(\log N)$.
@@
Now among the indices $i\in[s]$ such that $\beta_{i}$ is odd, let $i^{*}$ be the index such that $\beta_{i}$ is minimal. Then $\gamma=\beta_{i^{*}}$, and hence
@@
      • Otherwise, we can assume that assume that $B$ is squarefree and $M/B$ is square (with probability $1-\mathsf{negl}(\log N)$). Then by repeatedly dividing $M$ by $B$, we can find integers $k,\gamma$ such that $M=k\cdot B^{\gamma}$ and $k$ is not divisible by $B$.
@@
Algorithm 5.1 Fully factoring special integers (see Theorem 5.3)
@@
A special integer $N$ can be fully factored with success probability $1-\mathsf{negl}(\log N)$ using $\omega((\log N)^{3/2})$ calls to a quantum circuit of size $\widetilde{O}(n)$.
@@
The circuit in Corollary 3.2 only requires $\widetilde{O}(n)$ gates. Then the algorithm in 5.2 can be realized with $T=\omega(\log N)$ calls to the circuit in Corollary 3.2 (here, the choice of $T$ is specified by Theorem 5.3.) Finally, the algorithm in Theorem 5.3 can be realized with $O(\sqrt{\log N})$ calls to the algorithm of Lemma 5.2. Putting these together, the conclusion follows. ∎
@@ Acknowledgements.
The authors would like to thank Isaac Chuang, Mikhail Lukin, and Peter Shor for insightful discussions.

### v3

@@ 5 Completely Factoring Integers with Distinct Exponents in their Prime Factorization
We say that an integer $N$ is special if all the exponents in its prime factorization are distinct i.e. we can write $N=p_{1}^{\alpha_{1}}\ldots p_{r}^{\alpha_{r}}$ for distinct primes $p_{1},\ldots,p_{r}$ and distinct positive integers $\alpha_{1},\ldots,\alpha_{r}$.
Remark 6.
One might wonder whether special integers turn out to be classically easy to factor as well. We state some evidence that this is not the case here. Indeed, the density of these integers was studied by Aktaş and Murty [AM17], who showed that for any integer $N_{\mathrm{max}}$, the number of special integers $N\in[1,N_{\mathrm{max}}]$ is $\approx\frac{1.7N_{\mathrm{max}}}{\ln N_{\mathrm{max}}}$. The following classes of integers that are classically even slightly easier than general to completely factor do not contain enough elements in total to cover all special numbers:
@@
With this in mind, we now turn our attention to completely factoring special integers. The results and ideas in this section bear some high-level similarity to previous work by Yun [Yun76] in the context of factoring polynomials. Concretely, Yun shows that any polynomial $f(x)$ can be decomposed into a factorization $f(x)=g_{1}(x)^{\alpha_{1}}\ldots g_{r}(x)^{\alpha_{r}}$ where the $g_{i}$’s are squarefree and pairwise coprime and the $\alpha_{i}$’s are distinct. The starting point of Yun’s algorithm is the observation that if $f$ is divisible by the square of some polynomial $g$, then $g$ will divide $\gcd(f,f^{\prime})$. Similarly, in this section we will start from an algorithm for calculating squarefree decompositions of integers and obtain an algorithm for fully factoring special integers.
We first begin with a simple lemma. At a high level, we want to show that the $\Omega(1)$ success probability in Theorem 3.1 can be boosted to be very close to 1. This is not obvious since given $N=A^{2}B$ with $B$ squarefree, it may not be possible to efficiently determine whether the algorithm has succeeded in recovering $B$. We show that this is not difficult to work around.
@@
Let $N$ be a positive integer, with unique representation as $N=A^{2}B$ for $B$ squarefree. Moreover, we say that $N$ is very good if it is composite and neither squarefree nor a square.
@@
The below theorem and its proof bear some high-level similarity to a result by Yun [Yun76] that shows that a similar factorization can easily be carried out for polynomials.
@@
Then there exists another algorithm $\mathcal{B}$ that, given a special integer $N$ as input, recovers the complete prime factorization of $N$ with probability $1-\mathsf{negl}(\log N)$. This algorithm makes at most $O(\sqrt{\log N})$ calls to $\mathcal{A}$ with inputs $N^{\prime}$ that are always $\leq N$ and with repetition parameter $T=\omega(\log N)$. Outside of calls made to $\mathcal{A}$, the algorithm is classical and runs in time $\mathsf{poly}(\log N)$.
@@
Now among the indices $i\in[s]$ such that $\beta_{i}$ is odd, let $i^{*}$ be the index such that $\beta_{i}$ is minimal. Then $\gamma=\beta_{i^{*}}$ (where $\gamma$ is defined as computed in Algorithm 5.1), and hence
@@
      • Otherwise, we can assume that $B$ is squarefree and $M/B$ is square (with probability $1-\mathsf{negl}(\log N)$). Then by repeatedly dividing $M$ by $B$, we can find integers $k,\gamma$ such that $M=k\cdot B^{\gamma}$ and $k$ is not divisible by $B$.
@@
Algorithm 5.1 Completely factoring special integers (see Theorem 5.3)
@@
A special integer $N$ can be completely factored with success probability $1-\mathsf{negl}(\log N)$ using $\omega((\log N)^{3/2})$ calls to a quantum circuit of size $\widetilde{O}(n)$.
@@
The circuit in Corollary 3.2 only requires $\widetilde{O}(n)$ gates. Then the algorithm in Lemma 5.2 can be realized with $T=\omega(\log N)$ calls to the circuit in Corollary 3.2 (here, the choice of $T$ is specified by Theorem 5.3.) Finally, the algorithm in Theorem 5.3 can be realized with $O(\sqrt{\log N})$ calls to the algorithm of Lemma 5.2. Putting these together, the conclusion follows. ∎
@@ Acknowledgements.
The authors would like to thank Henry Corrigan-Gibbs for giving a stimulating talk at CRYPTO 2024 that inspired the beginning of this project, and for useful subsequent discussions. The authors would also like to thank Isaac Chuang, Antoine Joux, Mikhail Lukin, and Peter Shor for insightful discussions. The authors would also like to thank Daniel J. Bernstein, Martin Ekerå, Laura Lewis, Oded Regev, and anonymous reviewers for helpful comments and feedback on the manuscript.